# Fluctuations of local plastic strain in granular media

### v1

@@ V Conclusion
We have experimentally studied, using a setup combining optical analysis and DAWS, the fluctuations of strain in a granular medium subjected to oscillatory stress. A detailed analysis of the sources of variability suggests that the strain fluctuations are governed by the local scale, which dominates the effect of the strain gradient along the vertical. Both methods provide similar results, allowing us to propose that the observations made by measuring the strain fluctuations through imaging the grains at the wall are relevant in the bulk. We find that the variance of the local strain is proportional to its mean, as seen in other amorphous systems. Furthermore, the fluctuations of plastic strain arise from the interplay between diffusion and strain anticorrelation, with the latter being more pronounced in dense packings, thereby limiting erratic rearrangements due to the reduced free volume.
@@ Appendix A DAWS
Figure 9: (a) The black curve represents a typical signal transmitted through the granular suspension, while the red curve corresponds to the coherent part, obtained by averaging $10^{2}$ signals from different packings after multiple full rotations of the drum. The window used to compute the signal decorrelation is taken at $t_{s}$ (see text). (b) Average spectrum of the transmitted signal, with the filtering window (in cyan) centered at $1$ MHz ±$\pm 10\%$.
The coherent part (red signal in Figure 9 a)) is obtained by averaging one hundred signals transmitted through independent configurations, each acquired after several full rotations of the drum. This coherent wave is subtracted from raw signals, which are subsequently filtered to $1$ MHz ±$\pm 10\%$ (see Figure 9 b)).
When the drum oscillates below the angle of repose, the relative motion of the grains induces changes in the ultrasound signals. The coherent part of the signal depends on the average properties of the medium, such as the effective bulk modulus and density.
@@ Appendix B Fitting parameters for the global compaction
$\phi=\phi_{f}-\frac{\phi_{f}-\phi_{0}}{1+B\ln\left(1+\frac{t}{\tau}\right)},$
@@ Appendix C Quadratic strain from DAWS
Figure 11: a) Mean square strain from DAWS plotted versus the average shear strain $\langle\overline{\epsilon_{xy,m}}\rangle$ and average backward shear strain $\langle\overline{\epsilon_{xy,r}}\rangle$, corresponding to pairs $[a_{k},c_{k}]$ and $[c_{k},a_{k+1}]$, respectively. b) Fluctuations of plastic strain from DAWS as a function of maximum strain $|\langle\overline{\epsilon_{xy,m}}\rangle|$. Circles: position $P1$. Losange: position $P3$. Stars: position $P2$. Red, blue, green and violet represent measurements at $\theta_{m}=21.8^{\circ},19.8^{\circ},16.8^{\circ}$ and $10.8^{\circ}$ respectively. In b), only positions $P1$ and $P3$ are represented. For a) and b), the values of $\langle\overline{\epsilon_{xy,m}}\rangle$ and $\langle\overline{\epsilon_{xy,r}}\rangle$ used for building the abcissa are the same as in Figure 7 a) and b), and determined from optics.

### v2

@@ V Conclusion
We have experimentally studied, using a setup combining optical analysis and DAWS, the fluctuations of strain in a granular medium subjected to oscillatory stress. A detailed analysis of the sources of variability suggests that the effect of the gradient of the mean strain along the vertical is much smaller than the strain fluctuations at the scale of the grain $d$. This allows to compare both optical and DAWS methods which, providing similar results, suggest that the observations made by measuring the strain fluctuations through imaging the grains at the wall are relevant in the bulk. We find that the variance of the local strain is proportional to its mean, consistent with observations in discrete or partitioned systems where mutual exclusion is a key governing mechanism. Furthermore, the fluctuations of plastic strain depend solely on the mean, regardless of the compaction state or the applied load, and arise from the interplay between a memory term and strain anticorrelations. Quantifying scaling behaviors [5, 32] across broader strain ranges and multiple system sizes, as well as investigating the effect of polydispersity [33] on strain variability, represent relevant directions for future research.
@@ Acknowledgements
Author contribution statement
All authors contributed equally to the paper.
Data availability statement
The data that support the findings of this study can be obtained from the corresponding author upon reasonable request.
@@ Appendix A DAWS
Figure 9: (a) The black curve represents a typical signal transmitted through the granular suspension, while the red curve corresponds to the coherent part, obtained by averaging $10^{2}$ signals from different packings after multiple full rotations of the drum. The window used to compute the signal decorrelation is taken at $t_{s}$ (see text). (b) Average spectrum of the transmitted signal, with the filtering window (in cyan) centered at 11 MHz $\pm 10\%$.
The coherent part (red signal in Figure 9 a)) is obtained by averaging one hundred signals transmitted through independent configurations, each acquired after several full rotations of the drum. This coherent wave is subtracted from raw signals, which are subsequently filtered to $1$ MHz $\pm 10\%$ (see Figure 9 b)).
@@ Appendix B Fitting parameters for the global compaction
$\phi=\phi_{f}-\frac{\phi_{f}-\phi_{0}}{1+B\log\left(1+\frac{t}{\tau}\right)},$
@@ Appendix C Quadratic strain from DAWS
Figure 11: a) Mean square strain from DAWS plotted versus the average shear strain $\langle\overline{\epsilon_{xy,m}}\rangle$ and average backward shear strain $\langle\overline{\epsilon_{xy,r}}\rangle$, corresponding to pairs $a_{k},c_{k}]$ and $c_{k},a_{k+1}]$, respectively. b) Fluctuations of plastic strain from DAWS as a function of maximum strain $|\langle\overline{\epsilon_{xy,m}}\rangle|$. Circles: position $P1$. Losange: position $P3$. Stars: position $P2$. Red, blue, green and violet represent measurements at $\theta_{m}=21.8^{\circ},19.8^{\circ},16.8^{\circ}$ and $10.8^{\circ}$ respectively. c) Fluctuations of plastic strain from DAWS as a function of mean plastic strain. In b) and c), only positions $P1$ and $P3$ are represented. For a), b), and c) the values of $\langle\overline{\epsilon_{xy,m}}\rangle$ , $\langle\overline{\epsilon_{xy,r}}\rangle$ and $\langle\overline{\epsilon_{xy,p}}\rangle$ used for building the abscissa are the same as in Figure 7 a) b) and d).